lflores8876:

Lenny and Lisa have obtained a 30-year, fixed-rate mortgage for $675250 with a 7.25% interest rate. They purchased 3 points and their rate is now 6.875%. Factoring in the cost of points, when is the break-even point on their mortgage?

Ninjoy:

Let's go ahead and list basic information first (what we know) -Amount of Mortgage - $675,250 -3 discount points = 3% of mortgage value -Rate of interest without points - 7.25% -Length - 30 Years -# of points - 3 points -Rate of interest with points - 6.875% In this case the amount paid would= 675250*3%= $20,257.5 Given this much, calculate the monthly payments with and without points, then subtract the two and you will calculate the break even point using the below formula: BREAK EVEN= Cost of Discount points/ MONTHLY savings